OpenStudy (anonymous):

2. Today Carly’s bank account contains $82. Yesterday she made a withdrawal of $25. What was the bank account balance before Carly made the withdrawal? a. Write a subtraction equation that models the problem. b. Solve the equation. c. Answer the question.

OpenStudy (anonymous):

i will give best response ;)

OpenStudy (anonymous):

2.) x - 25=82 x=25+82 x=107

OpenStudy (anonymous):

is that what i should write down for all the steps?

OpenStudy (anonymous):

No all the steps are X = Cost of T-Shirt Y = Cost of Book X + Y = 22.95 12.25 + Y = 22.95 Y = 22.95 - 12.25 = 10.70 X = current bank account balance Y = previous bank account balance W = net withdrawal Y = X + W Y = 82 + 25 Y = 107

OpenStudy (anonymous):

so what should i write for my answer? i need all the steps to get full credit

OpenStudy (anonymous):

Y = 107

OpenStudy (anonymous):

yeah iknow that but i just dont need the answer i need the steps before that also

OpenStudy (anonymous):

This should be your answer X = Cost of T-Shirt Y = Cost of Book X + Y = 22.95 12.25 + Y = 22.95 Y = 22.95 - 12.25 = 10.70 X = current bank account balance Y = previous bank account balance W = net withdrawal Y = X + W Y = 82 + 25 Y = 107

OpenStudy (anonymous):

107